A mixture of gases contains 5.77 moles of oxygen (O2), 0.94 moles of hydrogen (H2), and 9.19 moles of nitrogen (N2). Calculate the partial pressures of the gases if the total pressure is 4.00 atm at a certain temperature.


1
Expert's answer
2021-12-03T08:32:02-0500

"5.77+0.94+9.19=15.9"


"P_{O_2}=\frac{5.77}{15.9}×4.00=1.45"


"P_{H_2}=\frac{0.94}{15.9}×4.00=0.24"


"P_{N_2}=\frac{9.19}{15.9}×4.00=2.31"